### v1

@@ 3.2.2   Link-level Analysis
Effective traffic management at the link level is essential for maintaining smooth traffic flow and ensuring safety on individual road links within transportation networks. This process includes accurate predictions and estimations of traffic conditions such as flow, speed, and travel time. DGMs play a pivotal role at this scale by providing advanced insights that support real-time traffic management, routing optimization, and congestion control. By leveraging the power of DGMs, traffic managers can respond more effectively to changing conditions on each link and anticipate future variations with greater precision. As discussed earlier, one of the benefits of using DGMs in estimation and prediction tasks is their ability to learn the characteristics and distributions from training data, enabling them to generate realistic and reliable data for effective traffic decision-making (Lin et al., 2018; Chen et al., 2018; Zhang et al., 2019a, b; Impedovo et al., 2019; Zang et al., 2019; Xu et al., 2020a; Li et al., 2020b; Yu et al., 2020a; Aibin et al., 2021; Sun et al., 2021b; Zhang et al., 2021a; Zhou et al., 2020; Song et al., 2021; Wang et al., 2022a; Zhao et al., 2022; Mo et al., 2022a; Xu et al., 2022; Khaled et al., 2022). For example, Liang et al. (2018) introduced a GAN-based model that incorporates two LSTM networks and traffic-flow theory for spatial-temporal traffic state estimation. The GAN generates realistic traffic-state values by learning from the underlying data distributions. Similarly, Li et al. (2019b) presented DeepGTT, a VAE-based model designed to predict travel time distributions by integrating real-time traffic data and spatial features. The VAE enables the model to generate realistic travel time distributions even under conditions of data sparsity and variability. In another study, Xu et al. (2020b) developed the Graph Embedding GAN (GE-GAN), which selects adjacent links to estimate road traffic speed and volume more accurately. To be more specific, it uses Graph Embedding (GE) techniques to represent the spatial characteristics of road networks, while a Wasserstein GAN (WGAN) is employed to generate traffic state data. Jin et al. (2022) proposed the PL-WGAN, integrating the advantages of multiple deep learning units, including WGAN, GCN, RNN, and attention mechanisms, for predicting short-term traffic speed in urban networks. The WGAN in this framework generates realistic traffic speed data by learning from historical patterns and control parameters. Yu et al. (2020b) introduced a framework combining clustering and generative models to predict taxi hotspots. This model, which integrates LSTM with CGAN, learns the spatiotemporal distribution of taxi hotspots to generate accurate predictions. Moreover, Zhou et al. (2020) proposed a Bayesian framework that combines VAEs with GNNs for robust traffic prediction. The model benefits from the generative capabilities of VAEs and Normalizing Flow to capture multimodal traffic data distributions, addressing the uncertainty and complexity of road sensor networks. Recently, Mo et al. (2022b) successfully combined normalizing flows, GANs, and physics-informed deep learning to generate realistic traffic data while quantifying uncertainty. The model uses the normalizing flow model as the generator within the GAN framework to explicitly estimate data likelihood. Additionally, physics-informed deep learning is employed to integrate traffic domain knowledge, ensuring that the model adheres to the laws of traffic flow in data estimation.
@@
While many studies using DGMs focus on learning complex data distributions to generate accurate and realistic outcomes, another critical aspect is the density estimation of outputs. This feature is less explored but crucial for providing a probabilistic view of potential outcomes, enhancing decision-making under uncertainty (Koochali et al., 2019; Arnelid et al., 2019; Rasul et al., 2020, 2021; Wen et al., 2023). For instance, Koochali et al. (2019) presented ForGAN, a probabilistic forecasting model that uses conditional GANs to predict future sensory data by learning the full conditional probability distribution from historical data. This approach employs KL Divergence to measure and refine the accuracy of predicted distributions, ensuring they closely match true data distributions. Rasul et al. (2020) developed an autoregressive deep learning model for multivariate probabilistic time series forecasting using conditioned normalizing flows. The model captures complex dependencies between multiple time series by modeling the joint distribution of future observations conditioned on past data, significantly improving prediction accuracy and uncertainty estimation. Rasul et al. (2021) introduced TimeGrad, which applies a denoising diffusion model to the same problem. TimeGrad samples from the data distribution at each time step by estimating its gradient and transforming noise into meaningful data through a Markov chain using Langevin sampling. Similarly, Wen et al. (2023) developed DiffSTG, a framework for probabilistic Spatio-Temporal Graph (STG) forecasting that combines Spatio-Temporal GNNs (ST-GNNs) with DDPMs. DiffSTG captures both spatial and temporal dependencies in STG data while modeling inherent uncertainties. This model addresses the inefficiencies of TimeGrad in long-term forecasting by using a non-autoregressive approach to predict multiple future horizons simultaneously. In another study, Feng et al. (2024) introduced the Latent Diffusion Transformer (LDT) for high-dimensional multivariate probabilistic time series forecasting. LDT employs a latent space approach combined with a diffusion-based conditional generator and a symmetric statistics-aware autoencoder to enhance the expressiveness and manageability of forecasting complex time series data. These studies demonstrate the robust capabilities of DGMs in link-level analysis, offering essential tools for predicting and managing traffic conditions on road networks, which further contributes to more efficient and responsive traffic management systems.
@@ 3.2.3   Region-level Analysis
At the regional level, traffic management includes estimating and predicting traffic conditions across large geographic areas encompassing multiple road networks and transportation systems. This type of analysis is essential for understanding broader traffic patterns and making informed decisions that impact urban planning, resource allocation, and emergency response strategies. Key tasks at the region level include predicting and estimating average speeds across the network, estimating origin-destination (OD) flows to understand travel patterns, forecasting traffic demand to anticipate future needs, analyzing overall traffic flow to identify congestion areas, and performing crash analysis to enhance network safety. The complexity of managing traffic at the regional level arises from the need to integrate data from diverse sources and understand the dynamic interactions between various components of the transportation network. From this perspective, DGMs are particularly well-suited due to their capability to learn from extensive datasets, capture complex spatial and temporal patterns, and model uncertainties of output results. This enables DGMs to provide accurate and robust estimations and predictions crucial for strategic decision-making and enhancing the efficiency and safety of transportation systems. Many studies have applied DGMs in the region-level traffic estimation and prediction (Saxena and Cao, 2019; Zhang et al., 2020b; Li et al., 2020a; Kakkavas et al., 2021; Wu et al., 2021; Feng et al., 2021; Kakkavas et al., 2021; Naji et al., 2021; Wang et al., 2022b; Li et al., 2022c, b; Lin et al., 2023b; Zhang et al., 2023a; Zarei et al., 2024; Li et al., 2024b; Shao et al., 2024). For example, Yu and Gu (2019) proposed a deep neural network architecture called the Graph Convolutional Generative Autoencoder (GCGA) to address real-time speed estimation problems. In GCGA, the Graph Convolutional Network (GCN) extracts spatial characteristics, and an autoencoder-based GAN uses these features to generate traffic speed maps from incomplete data. Yu et al. (2019) developed a deep generative model framework that integrates a modified DBSCAN algorithm and a CGAN model to predict taxi-passenger demand. The model accounts for spatial, temporal, and external dependencies of input data, providing accurate demand predictions. In another study, Wang et al. (2020b) introduced the SeqST-GAN, a sequence-to-sequence GAN model comprising LSTM and CNN for multi-step urban crowd flow prediction. This model enhances long-term prediction accuracy by treating crowd flow data as "image frames" and incorporating contextual features. Huang et al. (2022b) presented the Dynamic Multi-Graph Convolutional Network with GAN framework (DMGC-GAN) for predicting OD-based ride-hailing demand. DMGC-GAN effectively addresses high sparsity and complex spatio-temporal correlations in OD demand data by leveraging dynamic, directed OD graphs within a GAN framework. Furthermore, Lin et al. (2023b) developed Curb-GAN, a conditional GAN-based model for estimating urban traffic conditions under various travel demand scenarios. Curb-GAN integrates dynamic convolutional layers and self-attention mechanisms to capture spatial and temporal correlations, producing accurate and realistic traffic estimations. Yuan et al. (2023) introduced a model based on the DDPM for traffic matrix estimation. The DDPM learns the distribution of traffic matrices and improves estimation accuracy by incorporating a preprocessing module and framing the estimation task as a gradient-descent optimization problem. These studies illustrate how DGMs can effectively manage and predict traffic conditions at the regional level, providing valuable insights for strategic planning and operational decision-making across large transportation networks.
@@ 3.3 DGM for Unsupervised Representation Learning
Unsupervised representation learning involves training a model to learn useful features or representations of the data without requiring labeled inputs. This method is valuable in environments where labeled data is sparse or expensive, which is often the case in transportation datasets. DGMs, such as VAEs are adept at discovering intricate structures in unlabeled data by learning to compress data and reconstruct it back into the original space. In other words, DGMs learn to encode data into a compact, latent space—a lower-dimensional representation of the original data—which preserves much of the information but in a more compressed form. The latent vector is a compact representation of the input data that captures its most critical features. These vectors are the output of the model’s encoder component and serve as a compressed knowledge base of the data, containing the essential information needed to reconstruct or generate new data points. In transportation research, the manipulation and analysis of these latent vectors allow researchers to infer traffic patterns, and anomalies that might not be apparent in the high-dimensional original data.
@@
One primary research direction is feature extraction and classification. It gains popularity in the field of transportation when employing DGMs, particularly because of their potent capability to discern and categorize complex, multi-dimensional data without direct supervision. Many studies have used the latent space in various classification tasks such as transportation mode identification, driving behavior analysis, and anomaly detection (Krajewski et al., 2018; Rákos et al., 2021, 2020; Yao and Bekhor, 2022; Rakos et al., 2021; De Miguel et al., 2022; Xie et al., 2023; Santhosh et al., 2021; Islam et al., 2021; Krajewski et al., 2019; Ding et al., 2019; Yuan et al., 2023; Kim et al., 2021). For example, Gupta et al. (2018) proposed Social GAN and explored how variations in the latent variables of a generative model influence the generated human trajectories. They discovered that specific directions in the latent space correspond to changes in trajectory direction and speed, providing insights into controlling the types of behaviors the model can generate. Zhang et al. (2022c) introduced the Dirichlet VAE (DirVAE) to identify transportation modes from GPS trajectory data. The latent space of the DirVAE visualizes and classifies different transportation modes, demonstrating the utility of latent space in distinguishing between various modes of transportation. In another study, Boquet et al. (2020) used the VAE to represent complex, high-dimensional traffic data. The VAE learns a low-dimensional latent space representation that captures the underlying patterns and dependencies within the traffic data. This latent space helps visualize the distribution of different traffic data types, such as normal and anomalous patterns. Neumeier et al. (2021) proposed an unsupervised model based on the VAE architecture to predict vehicle trajectories with an interpretable latent space. The latent space can be analyzed to understand and predict lane-changing maneuvers, enhancing the model’s utility in studying vehicle behaviors. Chen et al. (2021c) developed a method for generating realistic traffic flow data using GANs with semantic latent code manipulation. By exploring and manipulating the semantic representations in the latent space, their model generates traffic flow data that accurately reflects realistic patterns and conditions. Moreover, Santhosh et al. (2021) proposed a novel approach combining CNN and VAE to classify vehicle trajectories and detect anomalies. This hybrid model addresses the challenges of classifying time-series data of varying lengths and identifying anomalies such as lane violations, sudden speed changes, and vehicles moving in the wrong direction. The visualization of the latent space in their model provides a clear understanding of how the VAE encodes trajectory data and distinguishes between normal and anomalous patterns. These studies highlight the effectiveness of DGMs in feature extraction and classification tasks within transportation. By fully utilizing the latent space, researchers can find intricate patterns and dependencies, facilitating more accurate and insightful analysis of transportation data.

### v3

@@ 3.2.2   Link-level Analysis
Effective traffic management at the link level is essential for maintaining smooth traffic flow and ensuring safety on individual road links within transportation networks. This process includes accurate predictions and estimations of traffic conditions such as flow, speed, and travel time. DGMs play a pivotal role at this scale by providing advanced insights that support real-time traffic management, routing optimization, and congestion control. By leveraging the power of DGMs, traffic managers can respond more effectively to changing conditions on each link and anticipate future variations with greater precision. Table 6 provides an overview of recent studies, classifying them based on model type, observation/prediction length, sampling time, prediction samples, dataset, and evaluation metrics.
Table 6: A summary of recent studies in link-level analysis using DGM
Author Model Task Input / Prediction Length Sampling Time Prediction Samples Dataset Evaluation Metrics
Yu and Gu (2019) GCGA (GAN-based) Speed Estimation - - 1 Cologne Speed Data MAPE
Li et al. (2019b) DeepGTT (VAE-based) Travel Time Distribution Estimation - - 1 Chinese Provincial Capital City Taxi Dataset RMSE, MAE
Xu et al. (2020b) GE-GAN Volume Estimation Speed Estimation - - 1 PEMS Seattle Dataset RMSE, MAE, MAPE
Yu et al. (2020b) LSTM-CGAN Taxi Hotspots Prediction 60min/10min 10min 1 Beijing Taxi trajectory Taximeter Data 1-Recall, FPR, ROC, Section Consistency
Zhou et al. (2020a) VGRAN (VAE + Flow) Speed Prediction 60min/5-60min 5min 1 METR, PEMS RMSE, MAE, MAPE
Rasul et al. (2020) Flow Flow Prediction Volume Prediction Flexible/24hr Flexible/12hr 1hr 0.5hr 100 PEMS-SF NYC Taxi Data CRPSsum, MSE
Rasul et al. (2021) TimeGrad (Diffusion-based) Flow Prediction Volume Prediction Flexible/24hr Flexible/12hr 1hr 0.5hr 100 PEMS-SF NYC Taxi Data CRPSsum
Wen et al. (2023) DiffSTG (Diffusion-based) Flow Prediction 1hr/1hr 5min 100 PEMS08 CRPS, RMSE, MAE
Feng et al. (2024) LDT (Diffusion-based) Flow Prediction Volume Prediction Flexible/24hr Flexible/12hr 1hr 0.5hr 100 PEMS-SF NYC Taxi Data CRPSsum, MSE
@@
As discussed earlier, one of the benefits of using DGMs in estimation and prediction tasks is their ability to learn the characteristics and distributions from training data, enabling them to generate realistic and reliable data for effective traffic decision-making (Lin et al., 2018; Chen et al., 2018; Zhang et al., 2019a, b; Impedovo et al., 2019; Zang et al., 2019; Xu et al., 2020a; Li et al., 2020b; Zhou et al., 2020a; Yu et al., 2020a; Aibin et al., 2021; Sun et al., 2021b; Zhang et al., 2021a; Song et al., 2021; Jin et al., 2022; Wang et al., 2022a; Zhao et al., 2022; Mo et al., 2022a; Xu et al., 2022; Khaled et al., 2022). For example, Yu and Gu (2019) proposed a deep neural network architecture called the Graph Convolutional Generative Autoencoder (GCGA) to address real-time speed estimation problems. In GCGA, the Graph Convolutional Network (GCN) extracts spatial characteristics, and an autoencoder-based GAN uses these features to generate traffic speed maps from incomplete data. Similarly, Li et al. (2019b) presented DeepGTT, a VAE-based model designed to predict travel time distributions by integrating real-time traffic data and spatial features. The VAE enables the model to generate realistic travel time distributions even under conditions of data sparsity and variability. In another study, Xu et al. (2020b) developed the Graph Embedding GAN (GE-GAN), which selects adjacent links to estimate road traffic speed and volume more accurately. To be more specific, it uses Graph Embedding (GE) techniques to represent the spatial characteristics of road networks, while a Wasserstein GAN (WGAN) is employed to generate traffic state data. Yu et al. (2020b) introduced a framework combining clustering and generative models to predict taxi hotspots. This model LSTM-CGAN, which integrates LSTM with CGAN, learns the spatiotemporal distribution of taxi hotspots to generate accurate predictions. Moreover, Zhou et al. (2020a) proposed a Bayesian framework that combines VAEs with GNNs for robust traffic prediction. The model benefits from the generative capabilities of VAEs and Normalizing Flow to capture multimodal traffic data distributions, addressing the uncertainty and complexity of road sensor networks.
@@
Many studies of DGMs focus on learning complex data distributions to generate accurate and realistic outcomes. However, an equally important but less explored aspect is the density estimation of outputs. Without probabilistic modeling, it is challenging to differentiate forecasts between low and high-noise scenarios, which is essential for offering a full probabilistic view and enhancing decision-making under uncertainty (Arnelid et al., 2019; Rasul et al., 2020, 2021; Wen et al., 2023; Lin et al., 2024). Rasul et al. (2020) introduced an autoregressive deep learning framework for multivariate probabilistic time series forecasting that leverages conditioned normalizing flow. Their model learns the joint distribution of future observations by conditioning on historical data, enabling it to capture complex dependencies among multiple time series. This approach not only improves prediction accuracy but also provides robust uncertainty estimates. Building on this work, Rasul et al. (2021) proposed TimeGrad, which employs a denoising diffusion model for the same forecasting task. TimeGrad generates forecasts by progressively transforming white noise into meaningful data through a learned Markov chain. At each time step, it estimates the gradient of the data distribution and uses Langevin sampling to refine the predictions, further improving performance and uncertainty quantification. Similarly, Wen et al. (2023) developed DiffSTG, a framework for probabilistic Spatio-Temporal Graph (STG) forecasting that combines Spatio-Temporal GNNs (ST-GNNs) with DDPMs. DiffSTG captures both spatial and temporal dependencies in STG data while modeling inherent uncertainties. This model addresses the inefficiencies of TimeGrad in long-term forecasting by using a non-autoregressive approach to predict multiple future horizons simultaneously. In another study, Feng et al. (2024) introduced the Latent Diffusion Transformer (LDT) for high-dimensional multivariate probabilistic time series forecasting. LDT employs a latent space approach combined with a diffusion-based conditional generator and a symmetric statistics-aware autoencoder to enhance the expressiveness and manageability of forecasting complex time series data. Building on previous studies, DGMs have demonstrated robust capabilities in link-level analysis, providing essential tools for predicting and managing traffic conditions on road networks and thereby contributing to more efficient and responsive traffic management systems.
One of the limitations is that, compared to agent-level studies, there is a notable gap in research on density estimation at the link level. Furthermore, while some probabilistic models excel at quantifying uncertainty, they often show less performance in point forecast accuracy relative to deterministic methods. Their reliance on variational inference can also lead to inaccurate posterior estimates when data is limited, underscoring a trade-off between capturing uncertainty and achieving precise forecasts.
@@ 3.2.3   Region-level Analysis
At the regional level, traffic management includes estimating and predicting traffic conditions across large geographic areas encompassing multiple road networks and transportation systems. This type of analysis is essential for understanding broader traffic patterns and making informed decisions that impact urban planning, resource allocation, and emergency response strategies. Key tasks at the region level include predicting and estimating average speeds across the network, estimating OD flows to understand travel patterns, forecasting traffic demand to anticipate future needs, analyzing overall traffic flow to identify congestion areas, and performing crash analysis to enhance network safety. The complexity of managing traffic at the regional level arises from the need to integrate data from diverse sources and understand the dynamic interactions between various components of the transportation network. From this perspective, DGMs are particularly well-suited due to their capability to learn from extensive datasets, capture complex spatial and temporal patterns, and model uncertainties of output results(Liang et al., 2018; Saxena and Cao, 2019; Zhang et al., 2020c; Li et al., 2020a; Kakkavas et al., 2021; Wu et al., 2021; Feng et al., 2021; Kakkavas et al., 2021; Naji et al., 2021; Wang et al., 2022b; Mo et al., 2022b; Li et al., 2022c, b; Yuan et al., 2023; Lin et al., 2023b; Zhang et al., 2023a; Zarei et al., 2024; Li et al., 2024b; Shao et al., 2024). In Table 7, we summarize recent studies by considering the model used, task, observation/prediction length, sampling time, dataset, and evaluation metrics.
Table 7: A summary of recent studies in region-level analysis using DGM
Author Model Task Observation / Prediction Length Sampling Time Dataset Evaluation Metrics
Yu et al. (2019) cGAN Taxi-Passenger Demand Prediction 60min/10min 10min Beijing Taxi Dataset, New York City Taxi Dataset MSE, MAE, MAPE
Wang et al. (2020b) SeqST-GAN Crowd Flow Prediction 24hr/1-20hr 1hr BikeNYC, TaxiNYC RMSE, MAE
Huang et al. (2022b) DMGC-GAN OD Ride-Hailing Demand Prediction 200min/20min 20min New York City Taxi, Limousine Commission Ride-Hailing Dataset RMSE, MAE, MAPE, $R^{2}$
Li et al. (2022a) ADST-GAN Crowd Flow Prediction 24h/1-4h 1h BikeNYC, TaxiNYC RMSE
Zhang et al. (2020b) Curb-GAN Speed Prediction Inflow Prediction 12hr/1-12hr 1h Shenzhen Speed Dataset Taxi Inflow Dataset RMSE, MAPE
Yu et al. (2019) integrated a modified DBSCAN algorithm with a CGAN model built on LSTM to predict taxi-passenger demand. The model accounts for spatial, temporal, and external dependencies of input data, thereby providing accurate demand predictions. In another study, Wang et al. (2020b) introduced the SeqST-GAN, a sequence-to-sequence GAN model comprising LSTM and CNN for multi-step urban crowd flow prediction. This model enhances long-term prediction accuracy by treating crowd flow data as "image frames" and incorporating contextual features such as weather, holidays, and point of interests. Huang et al. (2022b) presented the Dynamic Multi-Graph Convolutional Network with GAN framework (DMGC-GAN) for OD-based ride-hailing demand prediction. The framework builds dynamic, directed OD graphs—capturing geographic adjacency, mutual attraction, and mobility associations—to tackle data sparsity and complex spatio-temporal dependencies. These graphs are processed by a Temporal Multi-Graph Convolutional Network (TMGCN) with a GRU and integrated into a GAN framework to refine predictions. Moreover, Li et al. (2022a) proposed the Attentive Dual-Head Spatial-Temporal GAN (ADST-GAN) to predict crowd flows. The model integrates attentive temporal and spatial mechanisms by combining ConvLSTM, self-attention, and a dual-head discriminator within the GAN framework. These components capture complex spatial-temporal dependencies in crowd flow data and mitigate overfitting, ensuring that the synthetic data generated closely mimics real-world traffic patterns. Furthermore, Zhang et al. (2020b) developed Curb-GAN, a conditional GAN-based model for estimating urban traffic conditions under various travel demand scenarios. Curb-GAN integrates dynamic convolutional layers to capture local spatial correlations along irregular road networks and self-attention mechanisms to model temporal dependencies, producing accurate and realistic traffic estimations. Previous research has shown the effectiveness of DGMs in managing and predicting traffic conditions at the regional level, providing valuable insights for strategic planning and operational decision-making across large transportation networks. However, they also face several limitations. Many models depend heavily on high-quality, abundant training data, which may not always be available—especially when sudden or atypical demand scenarios occur. Additionally, many approaches rely on grid-based or clustering partitioning of urban areas, a method that can oversimplify and fail to capture the complex geometries of real-world road networks. Finally, despite the use of DGMs to improve output quality, there remains a risk of generating overly smooth or “blurry” predictions that obscure critical localized variations.
@@ 3.3 DGM for Unsupervised Representation Learning
Unsupervised representation learning involves training a model to learn useful features or representations of the data without requiring labeled inputs. This method is valuable in environments where labeled data is sparse or expensive, which is often the case in transportation datasets. DGMs, such as VAEs, are adept at discovering intricate structures in unlabeled data by learning to compress data and reconstruct it back into the original space. In other words, DGMs learn to encode data into a compact, latent space—a lower-dimensional representation of the original data—which preserves much of the information but in a more compressed form. The latent vector is a compact representation of the input data that captures its most critical features. These vectors are the output of the model’s encoder component and serve as a compressed knowledge base of the data, containing the essential information needed to reconstruct or generate new data points. In transportation research, the manipulation and analysis of these latent vectors allow researchers to infer traffic patterns, and anomalies that might not be apparent in the high-dimensional original data.
@@
One primary research direction is feature extraction and classification. It gains popularity in the field of transportation when employing DGMs, particularly because of their potent capability to discern and categorize complex, multi-dimensional data without direct supervision. Many studies have used the latent space in various classification tasks such as transportation mode identification, driving behavior analysis, and anomaly detection (Krajewski et al., 2018; Rákos et al., 2021, 2020; Yao and Bekhor, 2022; Rakos et al., 2021; De Miguel et al., 2022; Xie et al., 2023; Santhosh et al., 2021; Islam et al., 2021; Ding et al., 2019; Yuan et al., 2023; Kim et al., 2021). For clarity, Table 8 summarizes these recent studies, detailing the models used, latent space functions, and datasets. One notable study by Zhang et al. (2022c) introduced the Dirichlet VAE (DirVAE) to identify transportation modes from GPS trajectory data. The latent space of the DirVAE visualizes and classifies different transportation modes, demonstrating the utility of latent space in distinguishing between various modes of transportation. In another study, Boquet et al. (2020) used the VAE to represent complex, high-dimensional traffic data. The VAE learns a low-dimensional latent space representation that captures the underlying patterns and dependencies within the traffic data. This latent space can be used in various tasks such as missing-data imputation, anomaly detection, traffic forecasting, Likewise, Neumeier et al. (2021) proposed an unsupervised model based on the VAE architecture to predict vehicle trajectories with an interpretable latent space. The latent space can be analyzed to understand and predict lane-changing maneuvers, enhancing the model’s utility in studying vehicle behaviors. Moreover, Chen et al. (2021c) developed a method for generating realistic traffic flow data using GANs with semantic latent code manipulation. By exploring and manipulating the semantic representations in the latent space, their model generates traffic flow data that accurately reflects realistic patterns and conditions. Additionally, Santhosh et al. (2021) proposed a novel approach combining CNN and VAE to classify vehicle trajectories and detect anomalies. This hybrid model addresses the challenges of classifying time-series data of varying lengths and identifying anomalies such as lane violations, sudden speed changes, and vehicles moving in the wrong direction. The visualization of the latent space in their model provides a clear understanding of how the VAE encodes trajectory data and distinguishes between normal and anomalous patterns.
These studies demonstrate that DGMs are effective for feature extraction and classification in transportation. Leveraging the latent space allows researchers to reveal intricate patterns and dependencies, leading to more precise and insightful analysis of transportation data. Future work should aim to develop methods that efficiently extract and utilize the information in latent representations, thereby enhancing model robustness and interpretability and deepening our understanding of transportation phenomena.
Table 8: A summary of recent studies in unsupervised representation learning using DGM
Author Model Latent Space Function Dataset
Zhang et al. (2022c) DirVAE Transportation Mode Classification Geolife V1.3, OSMNX MTL Trajet 206,2017
Boquet et al. (2020) VAE Missing-data Imputation, Anomaly Detection, Traffic Forecasting PEMS, UKM1, UKM4
Neumeier et al. (2021) DVAE Understanding Lane-Changing Maneuvers HighD
Chen et al. (2021c) GAN Linear Interpolation, Manipulation Modifies Traffic-Flow Properties PEMS
Santhosh et al. (2021) CNN-VAE Distinguish Normal and Anomalous Patterns T15, QMUL, 4WAY datasets